# Bidiagonal matrix factorisations related to multiple orthogonal polynomials

### v1

@@ 3.1. Gauss-Borel factorisation and multiple orthogonality
To make this representation clearer, we show the illustrative example $r=2$. Then,
@@
The following result explains how the Gauss-Borel factorisation of the moment matrix can be used to obtain representations for the type II multiple orthogonal polynomials and type I linear forms on the step-line. A more general result for mixed multiple orthogonal polynomials can be found in [2, Prop. 3].
@@
Suppose that the moment matrix $\mathrm{M}$ in (3.1) admits a Gauss-Borel factorisation $\mathrm{M}=\mathrm{A}\mathrm{B}$, where $\mathrm{A}$ is a unit-lower-triangular matrix and $\mathrm{B}$ is a non-singular upper-triangular matrix. Then, all the multiple orthogonal polynomials on the step-line with respect to $\left(v_{1},\cdots,v_{r}\right)$ exist and are unique and the sequences of type II multiple orthogonal polynomials $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ and type I linear forms $\left(q_{n}\right)_{n\in\mathbb{N}}$ on the step-line are determined by
$\mathrm{P}(x)=\big{[}P_{0}(x),P_{1}(x),P_{2}(x),\cdots\big{]}^{\top}=\mathrm{A}^{-1}\,\chi(x)\quad\text{and}\quad q=\big{[}q_{0},q_{1},q_{2},\cdots\big{]}=\sum_{j=1}^{r}\chi_{j}^{\top}(x)\mathrm{B}^{-1}v_{j}.$ (3.6)
Proof.
Recall that the moment matrix $\mathrm{M}$ admitting the Gauss-Borel factorisation is equivalent to the existence and uniqueness of all multiple orthogonal polynomials on the step-line with respect to $\left(v_{1},\cdots,v_{r}\right)$
Using (3.3), the linearity of the functionals $v_{i}$, and the definition of the vector $\mathrm{P}(x)$ in (3.6), we find that
$\mathrm{B}=\sum_{j=1}^{r}v_{j}\left[\mathrm{A}^{-1}\chi(x)\chi_{j}^{\top}(x)\right]=\sum_{j=1}^{r}v_{j}\left[\mathrm{P}(x)\chi_{j}^{\top}(x)\right].$ (3.7)
As a result, the entries of $B=\left(b_{n,\ell}\right)_{n,\ell\in\mathbb{N}}$ satisfy
$b_{n,rk+j}=v_{j+1}\left[x^{k}\,P_{n}(x)\right]\quad\text{for all}k,n\in\mathbb{N}\text{ and }0\leq j\leq r-1.$ (3.8)
Taking into account that $\mathrm{B}$ is an upper-triangular matrix, (3.8) implies that $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ satisfies the type II multiple orthogonality conditions (2.4), with $N_{n}=b_{n,n}$ for all $n\in\mathbb{N}$. Furthermore, because $\mathrm{A}^{-1}$ is a unit-lower-triangular matrix, $P_{n}(x)$ is a monic polynomial of degree $n$ for all $n\in\mathbb{N}$. Therefore, we can state that $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ are the type II orthogonal polynomials on the step-line with respect to $\left(v_{1},\cdots,v_{r}\right)$.
We denote $\mathrm{B}^{-1}=\left(c_{i,j}\right)_{i,j\in\mathbb{N}}$. Recalling the definition of the vector $q$ in (3.6) and that $\mathrm{B}^{-1}$ is a lower-triangular matrix, we have
$q_{n}=\sum_{j=1}^{r}\left(\chi_{i}^{\top}(x)\mathrm{B}^{-1}\right)_{n}\,v_{j}=\sum_{j=1}^{r}\sum_{m=0}^{\left\lfloor\frac{n-j+1}{r}\right\rfloor}c_{rm+j-1,n}\,x^{m}\,v_{j}\quad\text{for all }n\in\mathbb{N}.$ (3.9)
More explicitly, taking $n=rk+i$ with $n\in\mathbb{N}$ and $10\leq i\leq r-$ 1, we can write
$q_{rk+i}=\sum_{j=1}^{i+1}\sum_{m=0}^{k}c_{rm+j-1,rk+i}\,x^{m}\,v_{j}+\sum_{j=i+2}^{r}\sum_{m=0}^{k-1}c_{rm+j-1,rk+i}\,x^{m}\,v_{j}=\sum_{j=1}^{r}C_{rk+i,j}(x)\,v_{j},$ (3.10)
where $C_{rk+i,1}(x),\cdots,C_{rk+i,r}(x)$ are polynomials such that
$\deg\big{(}C_{rk+i,j}(x)\big{)}\leq\begin{cases}\hfil k&\text{if }1\leq j\leq i+1,\\ k-1&\text{if }i+2\leq j\leq r.\end{cases}$ (3.11)
Furthermore, we can use (3.3) and the linearity of the functionals $v_{i}$ to find that
$\mathrm{A}=\sum_{j=1}^{r}v_{j}\left[\chi(x)\chi_{j}^{\top}(x)\mathrm{B}^{-1}\right].$ (3.12)
As a result, recalling the definition of the vector $q$ in (3.6), $\mathrm{A}=q\left[\chi(x)\right]=\big{(}q_{k}\left[x^{n}\right]\big{)}_{k,n\in\mathbb{N}}$. Hence, taking into account that $\mathrm{A}^{-1}$ is a unit-lower-triangular matrix, $\left(q_{k}\right)_{n\in\mathbb{N}}$ satisfies the type I multiple orthogonality conditions (2.5). Therefore, recalling (3.10)-(3.11), we can state that $\left(C_{n,1}(x),\cdots,C_{n,r}(x)\right)_{n\in\mathbb{N}}$ and $\left(q_{n}\right)_{n\in\mathbb{N}}$ are, respectively, the sequences of type I multiple orthogonal polynomials and type I linear forms on the step-line with respect to $\left(v_{1},\cdots,v_{r}\right)$. ∎
Definition 3.2 (Leading coefficient of a type I linear form).
For any $n\in\mathbb{N}$, we refer to the diagonal entry $c_{n,n}\neq 0$ of $\mathrm{B}^{-1}$ as the leading coefficient of the type I linear form $q_{n}$ and denote it by $\mathsf{lc}(q_{n})$.
Note that $c_{n,n}$ is the leading coefficient of the type I multiple orthogonal polynomial $C_{n,i+1}(x)$ if $n=rk+i$ with $k\in\mathbb{N}$ and $10\leq i\leq r-$ 1.
3.2. Recurrence relations for the multiple orthogonal polynomials
Consider the matrix $\Lambda$, which has only ones in the supradiagonal and zeros everywhere else, i.e.,
Observe that:
  1. $\bullet$
    $\Lambda\chi(x)=x\,\chi(x)$,
  2. $\bullet$
    $\Lambda\chi_{1}(x)=x\,\chi_{r}(x)$ and $\Lambda\chi_{j}(x)=\chi_{j-1}(x)$ for $2\leq j\leq r$,
  3. $\bullet$
    $\Lambda^{r}\chi_{j}(x)=x\,\chi_{j}(x)$ for $1\leq j\leq r$.
@@ 3. $\bullet$
Using these transformations, we can get the recurrence relations verified by the type II polynomials and by the type I linear forms from the Gauss-Borel factorisation $\mathrm{M}=\mathrm{A}\mathrm{B}$ of the moment matrix (3.1). A more general result for mixed multiple orthogonal polynomials can be found in [2, Prop. 13].
@@
Suppose that the moment matrix $\mathrm{M}$ in (3.1) admits the Gauss-Borel factorisation $\mathrm{M}=\mathrm{A}\mathrm{B}$, where $\mathrm{A}$ is a unit-lower-triangular matrix and $\mathrm{B}$ is a non-singular upper-triangular matrix. Then, the sequences of type II multiple orthogonal polynomials $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ and type I linear forms $\left(q_{n}\right)_{n\in\mathbb{N}}$ on the step-line with respect to $\left(v_{1},\cdots,v_{r}\right)$ satisfy the recurrence relations
@@
$\mathrm{H}=\mathrm{A}^{-1}\Lambda\mathrm{A}=\mathrm{B}\left(\Lambda^{\top}\right)^{r}\mathrm{B}^{-1}.$ (3.16)
Proof.
Recalling from (3.6) that $\mathrm{P}(x)=\mathrm{A}^{-1}\chi(x)$ and using the transformation $\Lambda\chi(x)=x\,\chi(x)$, we find that
$\mathrm{A}^{-1}\Lambda\mathrm{A}\mathrm{P}(x)=\mathrm{A}^{-1}\Lambda\chi(x)=x\,\mathrm{A}^{-1}\chi(x)=x\,\mathrm{P}(x).$ (3.17)
Similarly, recalling from (3.6) that $q=\sum\limits_{j=1}^{r}\chi_{j}^{\top}(x)\mathrm{B}^{-1}v_{j}$ and using the transformation $\Lambda^{r}\chi_{j}(x)=x\,\chi_{j}(x)$ for all $1\leq j\leq r$ and the linearity of the functionals $v_{j}$, we also obtain
$q\mathrm{B}\left(\Lambda^{\top}\right)^{r}\mathrm{B}^{-1}=\sum_{j=1}^{r}\chi_{j}^{\top}(x)\mathrm{B}^{-1}v_{j}\,\mathrm{B}\left(\Lambda^{r}\right)^{\top}\mathrm{B}^{-1}=\sum_{j=1}^{r}\big{(}\Lambda^{r}\,\chi_{j}(x)\big{)}^{\top}\mathrm{B}^{-1}\,v_{j}=\sum_{j=1}^{r}x\,\chi_{j}^{\top}(x)\mathrm{B}^{-1}\,v_{j}=x\,q.$ (3.18)
Next, we use (3.3) and the linearity of the functionals $v_{i}$ to derive that
[…]
As a result,
$\mathrm{A}^{-1}\Lambda\mathrm{A}=\mathrm{A}^{-1}\Lambda\mathrm{M}\mathrm{B}^{-1}=\mathrm{A}^{-1}\mathrm{M}\left(\Lambda^{\top}\right)^{r}\mathrm{B}^{-1}=\mathrm{B}\left(\Lambda^{\top}\right)^{r}\mathrm{B}^{-1}.$ (3.20)
Therefore, the recurrence matrices in (3.17) and (3.18) are equal. Hence, we can define $\mathrm{H}$ such that (3.16), for which $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ and $\left(q_{n}\right)_{n\in\mathbb{N}}$ satisfy the recurrence relations in (3.15).
Furthermore, because $\mathrm{A}$ and $\mathrm{B}$ are, respectively, lower- and upper-triangular matrices, $\mathrm{H}=\mathrm{A}^{-1}\Lambda\mathrm{A}$ implies that $\mathrm{H}$ is a unit-lower Hessenberg matrix and $\mathrm{H}=\mathrm{B}\left(\Lambda^{\top}\right)^{r}\mathrm{B}^{-1}$ implies that $\mathrm{H}$ has $r$ sub-diagonals. ∎

### v2

@@ 2.2. Gauss-Borel factorisation of the moment matrix and multiple orthogonality
To make this representation clearer, we present the illustrative example $r=2$. Then,
@@
The following result, which we do not prove here as it can be found, in more generality, in the context of mixed multiple orthogonal polynomials in [2, Prop. 3], shows how the Gauss-Borel factorisation of the moment matrix can be used to obtain explicit expressions for type II multiple orthogonal polynomials and type I linear forms along the step-line.
@@
Suppose that the moment matrix $\mathrm{M}$ in (2.11) admits a Gauss-Borel factorisation $\mathrm{M}=\mathrm{C}\mathrm{D}$, where $\mathrm{C}$ is a unit-lower-triangular matrix and $\mathrm{D}$ is a non-singular upper-triangular matrix. Then, all the multiple orthogonal polynomials on the step-line with respect to $\left(v_{1},\cdots,v_{r}\right)$ exist and are unique and the sequences of type II multiple orthogonal polynomials $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ and type I linear forms $\left(q_{n}\right)_{n\in\mathbb{N}}$ on the step-line are determined by
$\mathrm{P}(x)=\big[P_{0}(x),P_{1}(x),P_{2}(x),\cdots\big]^{\top}=\mathrm{C}^{-1}\,\chi(x)\quad\text{and}\quad q=\big[q_{0},q_{1},q_{2},\cdots\big]=\sum_{j=1}^{r}\chi_{j}^{\top}(x)\mathrm{D}^{-1}v_{j}.$ (2.16)
The latter result means that the entries of the inverses of the triangular matrices appearing in the Gauss-Borel factorisation $\mathrm{M}=\mathrm{C}\mathrm{D}$ are the coefficients of the type II and type I multiple orthogonal polynomials on the step-line with respect to $\left(v_{1},\cdots,v_{r}\right)$, which we denote by $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ and $\left(A_{n,1}(x),\cdots,A_{n,r}(x)\right)_{n\in\mathbb{N}}$, respectively. To make this statement clearer, we explicitly write these polynomials as
$P_{n}(x)=\sum_{k=0}^{n}p_{n}(k)\,x^{k}\quad\text{and}\quad A_{n,i}(x)=\sum_{\ell=0}^{n_{i}-1}a_{n,i}(\ell)\,x^{\ell},\quad\text{for }n\in\mathbb{N}\text{ and }1\leq i\leq r,$ (2.17)
where, for $n=rm+k$, with $m\in\mathbb{N}$ and $0\leq k\leq r-1$, we have $n_{i}=m+1$ if $i\leq k+1$ and $n_{i}=m$ if $i\geq k+2$. Then, $\mathrm{C}^{-1}$ and $\mathrm{D}^{-1}$ are triangular matrices such that:
  • $\left(\mathrm{C}^{-1}\right)_{n,k}=p_{n}(k)$ when $k\leq n$,
  • $\left(\mathrm{D}^{-1}\right)_{m,n}=a_{n,i}(\ell)$ when $m=r\ell+i-1\leq n$, with $\ell\in\mathbb{N}$ and $1\leq i\leq r$.
@@
We can get the recurrence relations verified by the type II polynomials and by the type I linear forms from the Gauss-Borel factorisation $\mathrm{M}=\mathrm{C}\mathrm{D}$ of the moment matrix (2.11). We omit the proof, as it follows directly from a more general result on mixed multiple orthogonal polynomials given in [2, Prop. 13].
@@
Suppose that the moment matrix $\mathrm{M}$ in (2.11) admits the Gauss-Borel factorisation $\mathrm{M}=\mathrm{C}\mathrm{D}$, where $\mathrm{C}$ is a unit-lower-triangular matrix and $\mathrm{D}$ is a non-singular upper-triangular matrix. Then, the sequences of type II multiple orthogonal polynomials $\left(P_{n}(x)\right)_{n\in\mathbb{N}}$ and type I linear forms $\left(q_{n}\right)_{n\in\mathbb{N}}$ on the step-line with respect to $\left(v_{1},\cdots,v_{r}\right)$ satisfy the recurrence relations
@@
$\mathrm{H}=\mathrm{C}^{-1}\Lambda\mathrm{C}=\mathrm{D}\left(\Lambda^{\top}\right)^{r}\mathrm{D}^{-1},$ (2.19)
where $\Lambda$ is the infinite matrix with ones in the supradiagonal and zeros everywhere else, i.e.,